OpenStudy (anonymous):

PLEASE HELP !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Brittany is 3 years older than twice her younger brother;s age. If Brittany's brother is 5 years old , how old is Brittany ? Write an algebraic expression for Brittany's age and solve the problem . Explain your answer.

OpenStudy (lgbasallote):

hmm let x = brother brittany = 3 + 2x can you do the rest?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

HELP

sam (.sam.):

3+B=2Y 3+B=2(5) B=7

OpenStudy (anonymous):

explain , PLEASE how you solved this stupid problem k?

sam (.sam.):

Is it correct?

OpenStudy (anonymous):

I don;t know because "lgbasallote" said that the beginning went off like this brittany = 3 + 2x

OpenStudy (lgbasallote):

@.Sam. why 3 + B = 2Y? 3 years older than twice brother's age...

OpenStudy (anonymous):

so ,who's right?

OpenStudy (anonymous):

lgbasallote I perhaps side with you

sam (.sam.):

Go with @lgbasallote :D, I suck at thinking skills ,lol

OpenStudy (anonymous):

can somebody thoroughly explain the answer with a clear explanation?

OpenStudy (lgbasallote):

hahaha lol =))) anyway the equation is 3 + 2x = brittany age..brother's age = 5 = x so substitute

OpenStudy (anonymous):

I said clearly

OpenStudy (anonymous):

so

OpenStudy (anonymous):

@lgbasallote please explain

OpenStudy (anonymous):

The answer should be 13, if britaany bro 1s and shes twice as much 10 and +3 =13

OpenStudy (anonymous):

(5x2)+3

OpenStudy (anonymous):

so , who's right?

OpenStudy (anonymous):

me ang lgba have same answer

OpenStudy (anonymous):

X=5 , 3+2x, 3 + (2x5)=13

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

No reason for attitude, i am being nice and havent told you off or anything have i?